OpenStudy (anonymous):

implicit differentiation (dy/dx) of sq. rt (x+y) = 1 + x^2y^2

OpenStudy (anonymous):

so left hand side derivative with respect to x: d (sqrt(x+y))/dx = 0.5 * (x+y)^(-0.5) * (1+dy/dx) right hand side derivative with respect to x: d(1+x^2y^2)/dx = 2xy^2 + 2y*(dy/dx)*x^2 agree ?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

implicit means that you supose that actualy y is a function of x, y(x), without actually nowing the actual expretion for it. Then differentiate the \(\sqrt {x+y} = 1 + x^2y^2\) keeping this in mind.